OpenStudy (anonymous):

Line WX contains (−1, 2) and (4, 12) Line YZ contains points (5, 8) and (−2, −6). Lines WX and YZ are

OpenStudy (anonymous):

perpendicular, because the slopes are the same parallel, because the product of the slopes is −1 perpendicular, because the product of the slopes is −1 parallel, because the slopes are the same

OpenStudy (anonymous):

y-12=2(x-4) = wx

OpenStudy (anonymous):

yz = y-8=2(x-5)

OpenStudy (anonymous):

just use y-y1 = m (x-x1) in which m means gradient

OpenStudy (amilapsn):

@CaseyCarns How to find the slope of a line?

OpenStudy (anonymous):

I know it's either B or C

OpenStudy (anonymous):

of course they are parallel because they have the same gradient (slope)

OpenStudy (anonymous):

So it's C

OpenStudy (anonymous):

product of the slopes becomes -1 if only they are perpendicular

OpenStudy (anonymous):

answer is D